OpenStudy (anonymous):

Can someone please help explain this? Thanks The circular bull’s-eye of an archery target has a diameter of 24 cm, which is surrounded by 4 concentric rings with a width of 12 cm each. Draw the target in a coordinate plane with center at the origin. Write the equations of the circles that form the boundaries of the different regions of the target.

OpenStudy (anonymous):

The origin is at (0,0). The bull's eye has a diameter of 24 inches. so treat each x as one inch. Therefore, the bull's eye will have the points (-24,0), (24,0)

OpenStudy (anonymous):

Add 12 to the 24 four times to get the 4 other x's. (The other four will be on negative/positive, but have the same value)

OpenStudy (whpalmer4):

not inches, cm...but I agree, use 1 unit = 1 cm

OpenStudy (anonymous):

Bull's eye: (-24,0) (24,0) Ring 1: (-36,0) (36,0) Ring 2: (-48,0) (48,0) Ring 3: (-60,0) (60,0) Ring 4: (-72,0) (72,0)

OpenStudy (anonymous):

*I ment treat each x as 1 cm :P

OpenStudy (whpalmer4):

Probably the easiest thing to do here is figure out the various radii that the circles will have. They'll all have the same center at (0,0), so the equations will look like \[(x-h)^2+(y-k)^2 = r^2\]but \((h,k)=(0,0)\) so that becomes\[x^2+y^2=r^2\]and you just need to plug in the various values of \(r\) for the different circles.

OpenStudy (anonymous):

thank you